A spring has a spring Constant, k=90N/m calculate the extension of the spring when 0.45J of work are done stretching it?


1
Expert's answer
2022-01-21T08:25:43-0500

Given:

"k=90\;\rm N/m"

"W=0.45\;\rm J"


The work done by a stretched spring

"W=\frac{kx^2}{2}"

Hence, the stretching

"x=\sqrt{2W/K}=\sqrt{2*0.45/90}=0.1\:\rm m"